# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 5. Open Challenges and Research Directions
Despite significant advances in privacy-preserving digital currencies, numerous open challenges remain. These challenges stem from the unique intersection of cryptographic capabilities, monetary policy requirements, and societal expectations of privacy and auditability. We highlight a few directions that merit further investigation.
@@ 5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability. As pointed out in different previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)), there exists several theoretical gaps related to the definition of unified privacy models and auditability. Most protocols informally claim anonymity, unlinkability, or confidentiality; however, formal and composable definitions that encompass transaction graph obfuscation, timing metadata, and value-hiding under active adversaries are lacking. We also point out that protocols combining multiple primitives often lack formal, compositional security proofs in frameworks like UC (Canetti, 2000) or IITM (Kuesters et al., 2013). Auditability mechanisms (e.g., view keys, trapdoor disclosures) are rarely formalized. There is a lack of frameworks to reason about selective disclosure, accountability guarantees, or conditional anonymity under adversarial inspectors.
@@ 5.2. Scalability and Efficiency
Many cryptographic protocols and, more broadly, PETs remain computationally demanding. This includes fully homomorphic encryption, ZKPs, and MPC. However, real-world cryptocurrencies and CBDCs must support high transaction throughput with low latency and high reliability. To bridge this gap, future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many of the systems analyzed, yet several fundamental challenges remain. General-purpose ZKP schemes (e.g., zk-SNARKs, Bulletproofs) face issues related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2191919https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Additionally, support for efficiently generating proofs that can verify other proofs—known as recursive proof composition—is still limited, even though this capability is crucial for scalable applications like off-chain rollups. Developing ZKPs that are not only succinct and recursive, but also universally verifiable (i.e., verifiable by anyone without special access), remains a technically challenging and active area of research. Finally, no current ZKP scheme offers both strong post-quantum guarantees202020ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions, highlighting the need for further advances in quantum-resistant proof systems.
@@ 5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is generally incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. Several preliminary approaches explore this trade-off, yet significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. For instance, while fully underexplored, witness encryption holds promise for enabling advanced privacy-preserving policies, such as verifiable time locks or conditional disclosures without trusted third parties (see Appendix A.3.5). Additionally, emerging requirements such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure call for more flexible and context-aware privacy controls. Designing PETs that can support such adaptive and policy-compliant behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
@@ 5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems fundamentally rely on the concept of anonymity sets. However, managing these sets in a secure, scalable, and adaptive manner remains an open challenge, especially in decentralized or evolving environments. Many existing systems, such as those based on ring signatures, operate with fixed-size anonymity sets or employ heuristic-based decoy selection. These approaches struggle to maintain robustness as usage patterns shift, and they are vulnerable to poisoning attacks, where adversaries inject tainted or known outputs into the anonymity set to degrade privacy. Developing cryptographically sound methods for dynamic anonymity set construction that grows or shrinks with system usage, while resisting active attacks, remains an unsolved problem. Beyond the anonymity set, adaptive privacy must also account for the dynamic nature of system participants, particularly in regulatory or issuance roles. In systems like CBDCs or regulated cryptocurrencies, authorities responsible for auditing, compliance verification, or even token issuance may change over time due to policy shifts, jurisdictional realignments, or organizational restructuring. Supporting the evolution of designated entities, such as auditors or currency issuers, without compromising user anonymity or transaction confidentiality is a significant technical challenge. Current systems typically assume static trust anchors or fixed multi-signature sets. However, introducing cryptographic mechanisms to dynamically reconfigure these roles (e.g., via threshold cryptography, distributed key generation, or policy-based encryption) requires careful balance between adaptability and privacy. Specifically, open issues relate to securely update the set of auditing entities, policies evolving over time, or support role rotation or revocation while preserving unlinkability, correctness, and forward secrecy.
@@ 5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity in cryptocurrencies is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into cryptocurrency protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be deanonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Joint modeling of privacy across the ledger layer and network layer is underexplored. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in cryptocurrency systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This is a highly active area of research, particularly in adversarial environments where well-funded surveillance actors may target both layers simultaneously.
@@ 6. Conclusion
Designing digital currencies that balance user privacy with institutional auditability remains one of the most intricate and impactful challenges in modern cryptographic and systems engineering. Over the past decade, cryptocurrencies and CBDC prototypes have pushed innovation across cryptographic protocols, PETs, and system architectures. Yet, despite substantial progress, the design space remains fragmented, and no universally accepted blueprint has emerged.
This survey offers a structured and technically grounded overview of the evolving landscape of privacy-preserving digital currencies. We systematize the literature across decentralized cryptocurrencies and CBDCs, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives—such as anonymity, confidentiality, and auditability—with concrete technical mechanisms and design models. Our analysis traces the evolution of privacy-preserving digital currencies across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital currencies transition from academic prototypes to operational systems, ensuring robust privacy alongside regulatory compliance will require continued technical innovation and interdisciplinary coordination. By providing a unified taxonomy and analysis framework, this survey aims to support future research and guide the design of privacy-preserving digital currency systems.
@@ Acknowledgements.
We thank Marco Benedetti, Giuseppe Galano, Sara Giammusso, Antonio Muci, and Fabiana Rossi for their insightful discussions in the early stages of this work and for their valuable feedback throughout its development.References
@@
- Bech and Garratt (2017) Morten L Bech and Rodney Garratt. 2017. Central bank cryptocurrencies. BIS Quarterly Review September (2017).
@@
- Bowe et al. (2020) Sean Bowe, Alessandro Chiesa, Matthew Green, Ian Miers, Pratyush Mishra, and Howard Wu. 2020. ZEXE: Enabling Decentralized Private Computation. In Proc. of IEEE S&P’20. 947–964.
@@
- Saleem et al. (2024) Hamza Saleem, Amir Ziashahabi, Muhammad Naveed, and Salman Avestimehr. 2024. Hawk: Accurate and Fast Privacy-Preserving Machine Learning Using Secure Lookup Table Computation. arXiv:2403.17296
@@
- Zyskind et al. (2015) Guy Zyskind, Oz Nathan, and Alex Pentland. 2015. Enigma: Decentralized Computation Platform with Guaranteed Privacy. arXiv:1506.03471
@@
A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability
@@ Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) a breakdown of publication venues for the works reviewed in this survey, and (ii) a summary of mixing protocols referenced in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in Section A.3, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital currencies.
@@ A.1. Most Popular Publication Venues
| Venue Type | # of Publications | Frequent Venues ( 2 works) | |||||||
@@ A.2. Reviewed Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys key privacy-preserving digital currencies, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The corresponding legend in Table LABEL:tab:sota_mixing_legend extends the notation introduced in Table 1. These mixing protocols are discussed in Section 4.3.5, and the table serves as a reference to facilitate comparison across existing approaches in the literature.
@@
| CoinShuffle++ (Ruffing et al., 2016) | 2016 | ✓ | ✓ (B) | ✓ | ✓212121https://github.com/decred/cspp | ✓ | ✓ | ✓ | DC-net | ||
@@
| Obscuro (Tran et al., 2018) | 2018 | ✓ (B) | ✓ | ✓222222https://github.com/BitObscuro/Obscuro | ✓ | ✓ | TEE | ||||
@@
| MixEth (Seres et al., 2019) | 2019 | ✓ | ✓ (E) | ✓ | ✓232323https://github.com/seresistvanandras/MixEth | ✓ | VS | ||||
@@
| BlindHub (Qin et al., 2023) | 2023 | ✓ (B) | ✓ | ✓242424https://github.com/blind-channel/blind-hub | ✓ | ✓ | BS* | ||||
@@
A.3. Emerging Cryptographic Tools for Future Digital Currencies
In addition to the cryptographic tools discussed in the main survey, several advanced primitives offer promising tools for future research in privacy-preserving digital currencies. Although not yet widely adopted in practical systems, these techniques may enable more flexible or robust privacy features in regulated or decentralized contexts.
@@ A.3.1. Anonymous Credentials
[…]
@@ A.3.3. Functional Encryption
[…]

### v2

@@ 5. Open Challenges and Research Directions
Despite significant advances in privacy-preserving digital payment systems, numerous open challenges remain. These challenges stem from the unique intersection of cryptographic capabilities, monetary policy requirements, and societal expectations of privacy and auditability. We highlight a few directions that merit further investigation.
@@ 5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability. As pointed out in different previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)), there exists several theoretical gaps related to the definition of unified privacy models and auditability. Most protocols informally claim anonymity, unlinkability, or confidentiality, but lack formal, composable definitions covering graph obfuscation, timing metadata, or value-hiding under active adversaries. Systems combining multiple primitives rarely have UC (Canetti, 2000) or IITM (Kuesters et al., 2013) proofs, and auditability features (e.g., view keys, trapdoors) are seldom formalized. Frameworks for reasoning about selective disclosure, accountability, and conditional anonymity remain missing.
@@ 5.2. Scalability and Efficiency
Many cryptographic protocols and PETs remain computationally demanding (e.g., homomorphic encryption, ZKPs, and MPC). However, real-world payment systems must support high transaction throughput with low latency and high reliability. To bridge this gap, future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many of the systems analyzed, yet several fundamental challenges remain. General-purpose ZKP schemes (e.g., zk-SNARKs, Bulletproofs) face issues related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2212121https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Finally, no current ZKP scheme offers both strong post-quantum guarantees222222ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions, highlighting the need for further advances in quantum-resistant proof systems.
@@ 5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. Several preliminary approaches explore this trade-off, yet significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. For instance, while fully underexplored, witness encryption holds promise for enabling advanced privacy-preserving policies, such as verifiable time locks or conditional disclosures without trusted third parties (see A.3.5). Additionally, emerging requirements such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure call for more flexible and context-aware privacy controls. Designing PETs that can support such adaptive and policy-compliant behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
@@ 5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems rely on anonymity sets, yet managing them securely and adaptively remains an open challenge, particularly in decentralized or evolving environments. Current schemes, such as fixed-size ring signatures or heuristic decoy selection, degrade as usage patterns shift and are vulnerable to poisoning attacks, where adversaries insert tainted outputs to weaken privacy. Cryptographically sound methods for dynamic, attack-resistant anonymity set construction are still lacking. Beyond anonymity sets, adaptive privacy must address the evolving nature of system participants, especially in regulatory or issuance roles. In CBDCs and regulated payment systems, auditors, compliance entities, or issuers may change due to policy or organizational shifts. Supporting this evolution without compromising anonymity or confidentiality is difficult: most systems assume static trust anchors or fixed multi-signature sets. Cryptographic tools such as threshold cryptography, distributed key generation, or policy-based encryption could enable dynamic reconfiguration, but balancing adaptability with unlinkability, correctness, and forward secrecy remains unresolved.
@@ 5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity in crypto-asset systems is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into privacy-preserving payment systems protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be deanonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Joint modeling of privacy across the ledger layer and network layer is underexplored. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in digital payment systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This is a highly active area of research, particularly in adversarial environments where well-funded surveillance actors may target both layers simultaneously.
@@ 6. Conclusion
Designing digital payment systems that balance user privacy with institutional auditability remains one of the most intricate and impactful challenges in modern cryptographic and systems engineering. Over the past decade, crypto-asset systems and CBDC prototypes have pushed innovation across cryptographic protocols, PETs, and system architectures. Yet, despite substantial progress, the design space remains fragmented, and no universally accepted blueprint has emerged.
This survey provides a structured and technically grounded overview of the evolving landscape of privacy-preserving digital payment systems. We systematize the literature across untraceable payments, decentralized ledger-based solutions, and CBDCs, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives—such as anonymity, confidentiality, and auditability—with concrete technical mechanisms and design models. Our analysis traces the evolution of privacy-preserving digital payment systems across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital payment technologies progress from academic prototypes to deployed systems, balancing strong privacy guarantees with regulatory compliance will require ongoing technical advances and interdisciplinary efforts. This survey provides a unified taxonomy to guide future research and the design of privacy-preserving payment systems of the future.
@@ Acknowledgements.
We thank Marco Benedetti, Claudia Biancotti, Giuseppe Galano, Sara Giammusso, Antonio Muci, and Fabiana Rossi for their insightful discussions in the early stages of this work and for their valuable feedback throughout its development.References
@@
- Andola et al. (2021) Nitish Andola, Raghav, Vijay Kumar Yadav, S. Venkatesan, and Shekhar Verma. 2021. Anonymity on blockchain based e-cash protocols—A survey. Comput. Sci. Rev. 40 (2021), 100394.
@@
- ECB (2018) ECB. 2018. What is bitcoin? https://www.ecb.europa.eu/ecb-and-you/explainers/tell-me/html/what-is-bitcoin.en.html.
@@
- Loporchio et al. (2023) Matteo Loporchio, Anna Bernasconi, Damiano Di Francesco Maesa, and Laura Ricci. 2023. A survey of set accumulators for blockchain systems. Comput. Sci. Rev. 49 (2023), 100570.
@@
A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability
@@ Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) a breakdown of publication venues for the works reviewed in this survey, and (ii) a summary of mixing protocols referenced in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in A.3, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital payment systems.
@@ A.1. Most Popular Publication Venues
| Venue Type | Number of Publications | Frequent Venues ( 2 works) | |||||||
@@ A.2. Reviewed Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys key privacy-preserving digital payment and crypto-assets systems, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The corresponding legend in Table LABEL:tab:sota_mixing_legend extends the notation introduced in Table 1. These mixing protocols are discussed in Section 4.3.5, and the table serves as a reference to facilitate comparison across existing approaches in the literature.
@@
| CoinShuffle++ (Ruffing et al., 2016) | 2016 | ✓ | ✓ (B) | ✓ | ✓232323https://github.com/decred/cspp | ✓ | ✓ | ✓ | DC-net | ||
@@
| Obscuro (Tran et al., 2018) | 2018 | ✓ (B) | ✓ | ✓242424https://github.com/BitObscuro/Obscuro | ✓ | ✓ | TEE | ||||
@@
| MixEth (Seres et al., 2019) | 2019 | ✓ | ✓ (E) | ✓ | ✓252525https://github.com/seresistvanandras/MixEth | ✓ | VS | ||||
@@
| BlindHub (Qin et al., 2023) | 2023 | ✓ (B) | ✓ | ✓262626https://github.com/blind-channel/blind-hub | ✓ | ✓ | BS* | ||||
@@
A.3. Emerging Cryptographic Primitives
In addition to the cryptographic tools discussed in the main survey, several advanced primitives offer promising tools for future research in privacy-preserving digital payment systems. Although not yet widely adopted in practical systems, these techniques may enable more flexible or robust privacy features in regulated or decentralized contexts.
@@ A.3.1. Anonymous Credentials
Systems like Idemix (Camenisch and Van Herreweghen, 2002) and Coconut (Sonnino et al., 2020) allow users to prove possession of attributes (e.g., age, residency, KYC compliance) without revealing their identity. These systems support selective disclosure, enabling users to present only the minimal information required for a given interaction; for example, proving age above without disclosing birthdate. This selective transparency could be very important for digital payment systems subject to regulatory constraints (e.g., AML, KYC), as users can demonstrate compliance without full deanonymization. Moreover, advanced anonymous credential schemes could support unlinkability, preventing different credential uses from being correlated. However, at the time of writing, efficiently revoking credentials (e.g., after misuse or expiration) while preserving privacy is nontrivial and typically requires accumulator-based techniques or dynamic revocation protocols. As digital payments increasingly intersect with real-world identity frameworks, anonymous credentials offer a promising path for balancing privacy and regulatory needs.
@@ A.3.3. Functional Encryption
Functional encryption (Boneh et al., 2010) allows specific functions to be computed on encrypted data without revealing the underlying inputs. In contrast to traditional public-key encryption, where decryption reveals the entire message, functional encryption supports function-restricted decryption keys. In the context of digital payment systems, this could enable selective disclosure, e.g., allowing regulators to verify policy compliance or transaction conditions without full access to transaction details. For instance, a regulator could decrypt only whether a transaction exceeds a policy-defined threshold without learning the actual value. Potential applications may include: (i) allowing a central bank to verify that a transaction is below a predefined limit, without revealing the amount; (ii) enabling eligibility verification for financial services based on encrypted KYC attributes; and (iii) constructing privacy-preserving tax collection mechanisms that expose only aggregate (or policy-relevant) outcomes. Despite its promise, functional encryption remains largely theoretical for high-complexity functions and faces efficiency and key management challenges in practical deployments.